OpenStudy (anonymous):

–x + 3y = 9 change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. than please describe how to graph using slope intercept form

OpenStudy (anonymous):

can you solve for y here? can you show me what slope intercept form looks like?

OpenStudy (anonymous):

y = mx + b 3y = x + 9 y = x/3 + 3

OpenStudy (anonymous):

try getting what chism got for yourself!

OpenStudy (anonymous):

How would i graph that?